Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report page - 20130802


Observers: Paul Vreeswijk, Annalisa De Cia and Cristina Barbarino
Support team: Ofer Yaron 
TAT on duty: Nancy Elias-Rosa & Emma Walker
TO: Ariel Sanchez

************************************************

Subrun.night: 2.1 
Instruments: EFOSC2

*** WEATHER CONDITIONS

Night  Humidity  Wind speed      Sky 
 
beg.    10-20%       5-12 m/s     clouds 
mid.    20-50%       8-12 m/s     clear
end.    10-30%       5-8   m/s     clouds 


*** AFTERNOON CALIBRATIONS

PESSTO_calib (dome flats and arcs)
PESSTO_bias
+ NTT automatic morning calibrations


*** LOG 

Target                  UT        Type    Filter/Grism      Slit       Seeing   Airmass    Grade        Comments 
                                       (Fol/Cl)                                                                 (A/B/C)

EG247                 03:10   STD     g11,16,13     1.0,1.5   1.2-1.6"  1.28     B/C      Asymmetric/double-peak traces, bad focus (because of unstable/poor seeing)            
LSQ13azj            04:20   C         g13                1.5          1.6"         1.07     B         M-dwarf
LSQ13bbb          04:26  C          g13                1.5          1.6"          1.04     B         SO galaxy 
LSQ13bdc          05:30  C          g13                1.5          1.4"          1.08      B        II
LSQ13bbh          06:00  C          g13                1.0          1.2"          1.10      B        Ia 
Feige110            06:30  STD      g11,16,13     1.0,1.5    1.2"           1.11      B      
LSQ13bhl           06:55 C           g13                1.0          1.2"           1.02      B       Ia  
LSQ13bgk          07:16 C           g13                1.5          2.0"           1.02      B       II. Seeing jump (acquisition process repeated)
CSS130614:233746+144237  07:40 F  UBVRi, g11,g16  1.5   1.6"  1.40      B


*** NOTES

* Dome opened between 03 and 09 UT (clear-sky window)
* Due to poor/unstable seeing conditions, at the beginning of the observing night we need to focus each image 
* Observation of LSQ13bah aborted since no clear transient was detected in the aquisision images  (~04 UT), needs better seeing
* PSNJ01364816+1545310 aborted due to clouds (~09 UT)
Comments